Assignment 1
As a member of Ghana’s economic planning team for the 2020 fiscal year, answer the under
listed questions (a & b) given that:
Y = C + I + G + X - M....................................................... (1)
C = 0.8Yd............................................................................ (2)
M = 0.2Yd........................................................................... (3)
Yd = (1-t)Y.......................................................................... (4)
Additional information: t = 0.2Y (tax rate), G = 400 (government spending), X = 280
(exports) and I = 300 (investment)
(a) Solve for the equilibrium level of national income
(b) Assume a marginal propensity to import (my) of 0.3 and solve for the new
equilibrium national income
1
Expert's answer
2020-05-21T10:02:19-0400

(a) Solve for the equilibrium level of national income



"Y = C + I + G + X - M"

But:


"C = 0.8Yd \\
M = 0.2Yd \\
Yd = (1-t)Y\\
t = 0.2\\
G = 400 \\
 X = 280 \\
I = 300"


Therefore



"Y = 0.8(1 - 0.2)Y + 300 + 400 + 280 - 0.2(1 - 0.2)Y\\

Y = 0.64Y + 980 - 0.16Y\\
 0.52Y = 980\\

Y^* = \dfrac{980}{0.52} = \color{red}{1,884.62}"

(b) Assume a marginal propensity to import (my) of 0.3 and solve for the new

equilibrium national income



"Y = 0.8(1 - 0.2)Y + 300 + 400 + 280 - 0.3(1 - 0.2)Y\\

Y = 0.64Y + 980 - 0.24Y\\
 0.6Y = 980

Y^* = \dfrac{980}{0.6} = \color{red}{1,633.33}"